OpenStudy (anonymous):

What is the solution set of |2x – 3| – 10 = –1? 6 or –3 6 or 3 6 or –6 No solution

OpenStudy (theviper):

6 or -3

OpenStudy (anonymous):

how'd u get it?

OpenStudy (anonymous):

absolute value means how far it is from 0 on the number line. so the absolute value of 5, is 5 because on number line, it is 5 units from 0. however, the absolute value of -5 is still 5 because it is still 5 units from 0. so first get the absolute value by it self by adding 10 on each side l2x-3l = 9 then since its an absolute value 2x-3 can be positive or negative 9. I9I = 9 or I -9 I = 9 so you have to set 2x-3 equal to 9 and -9. so 2x-3=9 or 2x-3=-9. 2x=12 or 2x=-6 x=6 or x=-3 therefore, x=6 and x=-3. sub in numbers to check

hartnn (hartnn):

u can directly plug in the numbers from options and check whether it satisfies the equation like let me do it for 6 |12-3|-10=-1 9-10=-1--->true so 6 is a solution. now can u find another number?

OpenStudy (anonymous):

Ya I get it..........Thanks all of u

OpenStudy (theviper):

|2x – 3| – 10 = –1 =>2x-3-10=-1 =>2x-13=-1 =>2x=12 =>x=12/2=6 |2x – 3| – 10 = –1 =>-(2x-3)-10=-1 =>-2x+3-10=-1 =>-2x-7=-1 =>-2x=6 =>x=6/-2=-3

OpenStudy (anonymous):

Thanks

OpenStudy (theviper):

:)